danielle and julia each improved their yards by planting daylilies and ornamental grass. They bought their supplies from the same store. Danielle spend $60 of 10 daylilies and 5 bunches or ornamental grass. Julia spent $27 on 5 daylilies and 2 bunches on ornamental grass. Whats the cost of one bunch of ornamental grass and the cost of one daylily?
Answer by Cromlix(4381)

Hi there,
Make daylilies = x
Make ornamental grass = y
Danielle spend $60 of 10 daylilies
and 5 bunches or ornamental grass.
10x + 5y = 60.......(1)
Julia spent $27 on 5 daylilies
and 2 bunches on ornamental grass.
5x + 2y = 27........(2)
Simultaneous equations:
10x + 5y = 60.......(1)
5x + 2y = 27........(2)
Multiply Eq(2) by 2
10x + 5y = 60.......(1)
10x + 4y = 54.......(2)
Subtract Eq(2) from Eq(1)
y = 6
Substitute y = 6 in Eq(1)
10x + 5y = 60.......(1)
10x + 5(6) = 60
10x + 30 = 60
10x = 60 - 30
10x = 30
x = 3
Daylilies cost $3
Ornamental grass = $6
Hope this helps :-)